26 5squared - 11 5 squared

Answer by checkley77(12844)

26 5squared - 11 5 squared
If you mean 26*5^5-11*5^5 then:
26*25-11*25=(26-11)25=15825=375 ans.
If you mean:
(26/5)^2-(11/5)^2 then:
676/25-121/25=(676-121)/25=555/25=111/5 or 22.2 ans.